A used car salesman buys a car for $2,500 and sells it for $4,327. he then buys a camper with 2/3 of his profit and sells the camper for twice what he paid for it. How much money did the salesman make during these transactions?
Answer by mbarugel(146)

He bought a car for 2,500 and sold it for 4,327. Therefore, the profit he made from the sale of this car was: 4327 - 2500 = $1,827.
Then he bought a camper for 2/3 of that profit. So he bought it for (2/3)1827 = 1,218, and sold it for twice that money; that is, sold it for 1218*2 = $2,436. The profit from this sale was 2436-1218 = $1,218.
We conclude that the salesman made a total of 1827+1218 = $3,045 during these transactions.
I hope this helps!
Get more answers at Online Math Answers.com!